What is an equation for the following:
The perimeter of a pool table is 30ft. The table is twice as long as it is wide. What is the length of the pool table?

Answer by edjones(8007)

Let L=length and w=width.
2L+2w=P
.
There is a system of 2 equations:
L=2w
2L+2w=30
.
2*2w+2w=30
4w+2w=30
6w=30
w=5
L=10
.
Ed